the sum of the two consecutive odd numbers is 52. what are the two odd numbers?
Answer by phillywily(18)

Take x as the first odd number, x+2 as the second odd number(because odd numbers are 2 numbers apart).
x+(x+2)=52
2x+2=52
2x=52-2
2x=50
x=50/2
x=25
x+2=45+2
x+2=27
Hence, the two odd numbers are 25 and 27 each.